OpenStudy (anonymous):

Find the exact positive value of c if the area between the graph of y=x^2-c^2 and the x -axis is 36. Does that mean x=36? but then how do i solve for c?

OpenStudy (amistre64):

the area under the curve is equal to 36

OpenStudy (amistre64):

c is a constant here that you are trying to determine

OpenStudy (amistre64):

in other words; what is the integral of y = x^2 + C such that it equals 36

OpenStudy (amistre64):

youd need to find the solution for: 108 = x (x+sqrt(3C)) (x-sqrt(3C))

OpenStudy (amistre64):

or simply put..... since C = c^2 108 = x (x - csqrt(3))(x + csqrt(c))

OpenStudy (amistre64):

if you can read thru the typos... i hope its helpful

OpenStudy (amistre64):

something like this

OpenStudy (anonymous):

i came out with the answer of 3 by\[\int\limits_{c}^{-c}x^2-c^2\]

OpenStudy (anonymous):

taking the antiderivative and getting x^3/3-c^2x

OpenStudy (amistre64):

make sure you integrate half of that area then multiply it by 2 to check yourself

OpenStudy (anonymous):

then substituting c for x

OpenStudy (anonymous):

why would you have to take half

OpenStudy (amistre64):

do we want the interval and the value of "c" in the equation to be the same?

OpenStudy (amistre64):

because the graph of y = x^2 -C is centered on the y axis; and if you take the area of that, you get "zero' regardless of where you put it :)

OpenStudy (amistre64):

both sides cancel each other out.... just a hazard of the trade really

OpenStudy (anonymous):

oh I got it Thanks for the help!

OpenStudy (amistre64):

youre welcome